OpenStudy (anonymous):

Find the most general antiderivative of the function. Use C for any needed constant. f(x) = x(2 - x)^2

OpenStudy (anonymous):

I would multiply it out first; then it's simply addition of anti-derivatives.

OpenStudy (anonymous):

\[x(2-x)^{2} = x(4-4x+x^2) = 4x - 4x^2 + x^3\]

OpenStudy (anonymous):

o so its 2x^2=4/3x^3+1/4x^4

OpenStudy (anonymous):

sorry (-) not =

OpenStudy (amistre64):

yep; expand it, distribute it, then integrate is and add a C

OpenStudy (anonymous):

o ya forgot about taht

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yup yu got it

OpenStudy (amistre64):

:)